Please show your work step by step and make sure the answers are correct because I have no chance to reanswer for this question
What is the temperature at the plane where the wood meets the Styrofoam? Express your answer using two significant figures Constants A carpenter builds an exterior house wall with a layer of wood 3.2 cm thick on the outside and a layer of Styrofoam insulation 2.4 cm thick on the inside wall surface. The wood has k-0080w/(m·K), and he Styrofoam has k 0.010 W/(m K). The interior surface temperature is 20.0 °C, and the exterior surface temperature is-12.0 C T- OC You may want to review (Pages 565-572) Submit Request Answer For related problem-solving tips and strategies, you may want to view a Video Tutor Solution of Conduction through two bars in contact Part B What is the rate of heat flow per square meter through this wall? Express your answer using two significant figures. W/
Answer #1

Part A

Heat flow through wood is equal to heat flow through styrofoam.

k1L2(T - T1) = k2L1 (T2 - T)

here 1 is for wood and 2 is for styrofoam.

T = k1L2T1 + k2L1T2 /  k1L2 + k2L1

T = 0.080*0.024*(-12) + 0.010*0.032*20 / 0.080*0.024 + 0.010*0.032

T = -0.02304 + 0.0064 / 0.00224

T = -7.428 0C

-------------------------------------------------------------------------------

Part B

Heat flow = 0.080 * (-7.428 - (-12)) / 0.032

Heat flow = 11.43 W/m2